### v2

@@ B.1 Proof of Theorem 3
$|L_{B}(\mathcal{E},\Sigma,\gamma)|=2\{z_{1-\eta/2}\,\alpha(\mathcal{E},\Sigma,\gamma)+B\,\sqrt{\beta(\mathcal{E},\gamma)}\}.$
We can now follow verbatim the proof of Theorem 2, with $\delta(t):=\inf_{\mathcal{E},\Sigma,\gamma}2\{z_{1-\eta/2}\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)}+t\sqrt{\beta(\mathcal{E},\gamma)}\}$ with limits
$\delta(0)=2z_{1-\eta/2}\sqrt{\alpha^{\star}},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=2\sqrt{\beta^{\star}}.$
Taking the ratios and limits as in Step 3 of the proof of Theorem 2 the proof completes.
B.2 Proof of Theorem 4
It is easy to show that
$\check{R}_{B}\left(\mathcal{E},\Sigma,\gamma\right)\equiv\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)=\check{\beta}(\mathcal{E},\gamma)B^{2}+\check{\alpha}(\mathcal{E},\Sigma,\gamma),$
with $\check{\beta},\check{\alpha}$ as defined in Equation (12), where $\check{\beta}\geq 0$ by construction. In addition $\check{\alpha}>0$ and bounded away from zero by Assumption 1. We can therefore follow verbatim the proof of Theorem 1.
Appendix C Examples
C.1 Two parameters examples in Equation (3)
@@
C.2 Numerical example
Here, we present a numerical example in Setting 2.
@@ C.2 Numerical example
Figure 2 (top row) reports $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ evaluated at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$. In the bottom row, we plot the corresponding $\gamma_{2}^{\star}$.
@@
$\boldsymbol{\omega_{2}>\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Interior Increase in $\omega_{2}$
@@
Observation 1: The regret for $j=2$ decreases rapidly in $\omega_{2}$ and then decreases more slowly. In the first column, we vary $\omega_{2}$ with the first experiment $j=1$ having a smaller experimental variance, $v_{1}=v_{2}/2$. The panel shows that a small $\omega_{2}$ makes the regret of choosing experiment $2$ larger than that of choosing experiment $1$. This is because a small $\omega_{2}$ corresponds to a small bias from not choosing $j=2$. As $\omega_{2}$ increases, the max–regret curve for $j=2$ declines until it reaches $\omega_{2}\approx 1.25$; after this point, the regret curve for $j=2$ rises slowly, since a further increase in $\omega_{2}$ is associated with an increase in estimator variance.
@@
Figure 11: Regret comparisons. Top row: $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Bottom row: optimal weights $\gamma_{j}^{\star}$ for $j=1,2$, which explain how sensitivity and precision interact to drive the design switch at $x^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Appendix D Additional figures and tables
Table 5: Optimal shrinkage and experiment choice across different regimes in the presence of two parameters (Setting 2) The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases). The other cases correspond to boundary solutions.
@@ Appendix D Additional figures and tables
Figure 13: Optimal shrinkage $\gamma_{j}^{\star}$ by area as a function of total treated villages $n_{1}$ (from ten to fifty-two villages), under different constraints $E$. For $E<4$, the solution is interior ($0<\gamma_{j}^{\star}<1$) and increases with $n_{1}$, approaching one as experimental noise falls.

### v3

@@ A.2 Proof of Theorem 3
$\ell_{b}(W,\Sigma)=\underline{\tau}(\hat{\theta}(W))-\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b-z_{1-\eta/2}\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)},$
so
$\inf_{\|b\|_{l}\leq B}\ell_{b}(W,\Sigma)=\underline{\tau}(\hat{\theta}(W))-z_{1-\eta/2}\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}-\sup_{\|b\|_{l}\leq B}\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b.$
Since $b_{j}=0$ for $j\in\mathcal{I}$ and $\|b_{\mathcal{I}^{c}}\|_{l}\leq B$, the dual norm gives
$\sup_{\|b\|_{l}\leq B}\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b=B\,\sqrt{\beta_{l,\underline{\omega}}(W)}.$
Hence
$\inf_{\|b\|_{l}\leq B}\ell_{b}(W,\Sigma)=\underline{\tau}(\hat{\theta}(W))-z_{1-\eta/2}\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}-B\,\sqrt{\beta_{l,\underline{\omega}}(W)}.$
Similarly, for the upper bound,
$\sup_{\|b\|_{l}\leq B}u_{b}(W,\Sigma)=\bar{\tau}(\hat{\theta}(W))+z_{1-\eta/2}\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+B\,\sqrt{\beta_{l,\bar{\omega}}(W)}.$
Therefore the worst-case interval has length
$|L_{l,B}(W,\Sigma)|=\big(\bar{\tau}(\hat{\theta}(W))-\underline{\tau}(\hat{\theta}(W))\big)+z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big)+B\big(\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}\big).$
Let $b_{0}$ denote the true bias, with $\|b_{0}\|_{l}\leq B$. Under (17),
$\bar{\tau}(\hat{\theta}(W))-\underline{\tau}(\hat{\theta}(W))=(\bar{\omega}-\underline{\omega})^{\top}\theta+(\bar{\omega}-\underline{\omega})^{\top}\Gamma_{\Lambda}(W)\bar{g}_{\Sigma},$
so
$\mathbb{E}_{W,\Sigma,b_{0}}\big[\bar{\tau}(\hat{\theta}(W))-\underline{\tau}(\hat{\theta}(W))\big]=(\bar{\omega}-\underline{\omega})^{\top}\theta+(\bar{\omega}-\underline{\omega})^{\top}\Gamma_{\Lambda}(W)b_{0}.$
Subtracting the identified set length $\bar{\tau}(\theta)-\underline{\tau}(\theta)=(\bar{\omega}-\underline{\omega})^{\top}\theta$, we obtain
$\mathbb{E}_{W,\Sigma,b_{0}}\big[|L_{l,B}(W,\Sigma)|\big]-\big(\bar{\tau}(\theta)-\underline{\tau}(\theta)\big)=$
$(\bar{\omega}-\underline{\omega})^{\top}\Gamma_{\Lambda}(W)b_{0}+z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big)+B\big(\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}\big).$
Taking the worst case over $\|b_{0}\|_{l}\leq B$ implies
$\mathcal{L}_{l,B}(W,\Sigma)=z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big)+B\big(\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}+\sqrt{\beta_{l,\bar{\omega}-\underline{\omega}}(W)}\big),$
which yields the linear form $\mathcal{L}_{l,B}(W,\Sigma)=A(W,\Sigma)+B\,C_{l}(W)$.
Step 2 (oracle envelope and basic properties). Define
$\delta(t)~\equiv~\inf_{(W,\Sigma)\in\mathcal{D}^{\prime}}\{A(W,\Sigma)+t\,C_{l}(W)\},\qquad t\geq 0.$
By Assumption 3 (because either or both $\bar{\omega},\underline{\omega}$ are not zero vectors), $A(W,\Sigma)$ and $C_{l}(W)$ are bounded from above and $A(W,\Sigma)>0$ for all $(W,\Sigma)\in\mathcal{D}^{\prime}$, so $\delta(t)$ is finite and strictly positive for all $t\geq 0$. Being the pointwise infimum of affine functions of $t$, $\delta$ is concave and nondecreasing on $0,\infty)$. Moreover,
$\delta(0)=A^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=C_{l}^{\star}.$
Step 3 (quasi-convexity and boundary maximization). Fix $(W,\Sigma)$ and set $A\equiv A(W,\Sigma)$ and $C\equiv C_{l}(W)$. For $t\geq 0$ define
$\phi(t)~\equiv~\frac{A+t\,C}{\delta(t)}.$
The numerator is affine in $t$ and the denominator is positive and concave, so $\phi$ is quasi-convex on $0,\infty)$. Hence, on any compact interval $[0,T]$, $\max_{t\in[0,T]}\phi(t)$ is attained at the boundary $\{0,T\}$. Letting $T\to\infty$ and using the limits in Step 2,
$\sup_{t\geq 0}\phi(t)=\max\left\{\frac{A}{\delta(0)},\;\lim_{t\to\infty}\frac{A+t\,C}{\delta(t)}\right\}=\max\left\{\frac{A(W,\Sigma)}{A^{\star}},\;\frac{C_{l}(W)}{C_{l}^{\star}}\right\}.$
By the definition of $\tilde{\mathcal{R}}_{l}(W,\Sigma)$ and the identity $\mathcal{L}_{l,B}(W,\Sigma)=A(W,\Sigma)+BC_{l}(W)$, this yields
$\tilde{\mathcal{R}}_{l}(W,\Sigma)=\sup_{B\geq 0}\frac{\mathcal{L}_{l,B}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\mathcal{L}_{l,B}(W^{\prime},\Sigma^{\prime})}=\max\left\{\frac{A(W,\Sigma)}{A^{\star}},\;\frac{C_{l}(W)}{C_{l}^{\star}}\right\},$
which proves the theorem. ∎
A.3 Proof of Corollary 1
Whenever $\bar{\omega}=\underline{\omega}=\omega$, we can write $A(W,\Sigma)=\alpha_{\omega}(W,\Sigma)$ and $C_{l}(W)=\beta_{l,\omega}(W)$. From Theorem 3
[…]
Appendix B Examples
B.1 Two parameters examples in Equation (6)
@@
B.2 Numerical example
Here, we expand our discussion around Example 5.
Figure 9: Example for $j=2$ (regret). Different colors correspond to the regret for choosing either experiment. Figure reports $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Columns vary, respectively, (a) $\omega_{2}$ with $v_{2}=0.5$, $v_{1}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Consider choosing $\gamma_{2}^{\star}$ when experiment $j=2$ is more precise as $j=1$, with $\sigma_{1}^{2}=\sigma_{2}^{2}=1$, $v_{1}=1$, $v_{2}=0.5$, and $\omega_{1}=1$. The optimal $\gamma_{2}^{\star}$ depends not only on the features of arm $2$ but also on those of the alternative arm 2-2, $(v_{-j}^{2},\sigma_{-j}^{2},\omega_{-j})$, through the oracle benchmarks $(\alpha^{\star},\beta^{\star})$. Figure 8 illustrates this by plotting $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ at the optimal weight (top row) and the corresponding $\gamma_{2}^{\star}$ (bottom row), as we vary $\omega_{2}$, $v_{2}$, and $\sigma_{2}$ column by column. Varying $\omega_{2}$ yields two key effects: both very small and very large $\omega_{2}$ push $\gamma_{2}^{\star}$ toward one, but for different reasons. When $\omega_{2}$ is small, the oracle bias declines faster than the experimental variance (since $\beta^{\star}=\omega_{2}$), while when $\omega_{2}$ is large, the oracle bias is $\beta^{\star}=\omega_{1}$ and $\omega_{2}$ affects only the feasible experiment’s bias. Consequently, both regret and the optimal estimator depend on the oracle’s choice of which experiment it would run.
In the interior region for $\omega_{2}$, the optimal weight first moves toward the variance-optimal solution (as the oracle bias falls) and then rises again toward a bias-preferred solution. When we vary the experimental and observational variances, we obtain the expected comparative statics: $\gamma_{2}^{\star}$ decreases with the experimental variance and increases with the observational variance. These patterns translate into the regret of each arm (Figure 9). For arm $j=2$, regret initially falls sharply as $\omega_{2}$ rises and then declines more gradually once the bias component is already small. For arm $j=1$, regret is at best slightly decreasing for small changes in $\omega_{2}$ (when variance effects dominate) but eventually increases with $\omega_{2}$, because the oracle increasingly favors arm $2$ and the opportunity cost of selecting arm $1$ rises. Table 4 summarizes these patterns. We now provide further details.
@@ B.2 Numerical example
Figure 8 (top row) reports $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ evaluated at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$. In the bottom row, we plot $\gamma_{2}^{\star}$.
@@
$\boldsymbol{\omega_{2}>\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{1}^{2}\sigma_{1}^{2}\;+\;\omega_{2}^{2}\dfrac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ Interior Increase in $\omega_{2}$
@@
Observation 1: The regret for $j=2$ decreases rapidly in $\omega_{2}$ and then stabilizes/slowly increases. In the first column, we vary $\omega_{2}$ with the first experiment $j=1$ having a smaller experimental variance, $v_{1}=v_{2}/2$. The panel shows that a small $\omega_{2}$ makes the regret of choosing experiment $2$ larger than that of choosing experiment $1$. This is because a small $\omega_{2}$ corresponds to a small bias from not choosing $j=2$. As $\omega_{2}$ increases, the max–regret curve for $j=2$ declines until it reaches $\omega_{2}\approx 1.25$; after this point, the regret curve for $j=2$ rises slowly, since a further increase in $\omega_{2}$ is associated with an increase in estimator variance.
@@
Figure 10: Regret comparisons. Top row: $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Bottom row: optimal weights $\gamma_{j}^{\star}$ for $j=1,2$, which explain how sensitivity and precision interact to drive the design switch at $x^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{2}=0.5$, $v_{1}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Table 5: Optimal shrinkage and experiment choice across different regimes in the presence of two parameters (Example 5) The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases). The other cases correspond to boundary solutions.
@@
Appendix C Additional figures and tables
Figure 12: Experimental design for General Equilibrium Effects: Variance and bias regret of the proposed optimal design relative to the corresponding variance and bias oracles, as a function of total experimental sample size $n_{\text{tot}}$. The left panel considers the case with one feasible experiment and the right panel the case with two feasible experiments; values close to one indicate little loss relative to the oracle.
Figure 13: Experimental Design for General Equilibrium Effects: Optimal shrinkage weights $\gamma_{j}$ for each arm (UCT, CCT, and Training) as a function of the total experimental sample size $n_{\text{tot}}$. The left panel considers designs with one feasible experiment, while the right panel allows two feasible experiments, illustrating how the procedure gradually shifts weight from UCT toward the job program as sample size increases.
Figure 14: Site selection for external validity: Optimal shrinkage weight $\gamma_{j}^{\star}$ as a function of the treated villages in the experiment.
Figure 15: Site selection for external validity: Variance, bias, and adaptive regret of the Neyman (variance–optimal) allocation relative to the proposed optimal design, as a function of the total experimental sample size $n_{\text{tot}}$. Left panel: one feasible experiment; right panel: two feasible experiments.
Appendix D Review of adaptation regret with fixed design
This section reviews the definition of adaptation regret with a fixed design in Armstrong et al. (2024) and Tsybakov (1998) and illustrates distinctions from our framework.
Single parameter with a fixed design
Different from this paper, Armstrong et al. (2024) focus on a single parameter with two estimators under a fixed design where one of the two estimators is unbiased for the target estimand. In our notation, this corresponds to fixing a design $(E,\Sigma(E))$ with $E=\{1\}$, so there is an observational estimate $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}$ and an experimental estimate $\tilde{\theta}^{\exp}\in\mathbb{R}$ satisfying
$\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]=\theta_{1}+b,\qquad\mathbb{E}[\tilde{\theta}^{\exp}]=\theta_{1},\quad\Delta\tilde{\theta}=\tilde{\theta}^{\mathrm{obs}}-\tilde{\theta}^{\exp}$
for an unknown scalar bias $b\in\mathbb{R}$. Without loss, standardize each estimator as well as the estimand $\theta_{1}$ by the variance of the difference $\Delta\tilde{\theta}$. With this normalization, the authors further assume
$\begin{pmatrix}\tilde{\theta}^{\exp}\\ \Delta\tilde{\theta}\end{pmatrix}\sim\mathcal{N}\!\left(\begin{pmatrix}\theta_{1}\\ b\end{pmatrix},\,\Sigma\right),\quad\Sigma=\begin{pmatrix}\sigma_{\exp}^{2}&\rho\,\sigma_{\exp}\\[4.0pt] \rho\,\sigma_{\exp}&1\end{pmatrix}$,
where $\sigma_{\exp}^{2}$ is the variance of $\tilde{\theta}^{\exp}$, and $\rho$ is the correlation between the experimental estimate and $\Delta\tilde{\theta}$. The question of interest in Armstrong et al. (2024) is how to optimally combine $\tilde{\theta}^{\mathrm{exp}}$ and $\Delta\tilde{\theta}$ in a single estimator.
Non linear estimator and corresponding adaptation regret
Using invariance arguments, Armstrong et al. (2024) study estimators of the form
$\hat{\tau}_{\delta}=\tilde{\theta}^{\exp}+\rho\,\sigma_{\exp}\,\big\{\delta(\Delta\tilde{\theta})-\Delta\tilde{\theta}\big\}$,
for some measurable function $\delta:\mathbb{R}\to\mathbb{R}$ to be optimized. The authors show that for this choice, the adaptation regret worst case for $|b|\leq B$, can be written as
$A(B,\delta)=\frac{\displaystyle\mathrm{MSE}_{B}(\delta)}{R^{*}(B)},\qquad R^{*}(B)=\sigma_{\exp}^{2}\Big[(1-\rho^{2})+\rho^{2}\,r_{\mathrm{BNM}}(B)\Big]$
where $r_{\mathrm{BNM}}(\cdot)$ denotes the B-minimax risk in the bounded normal-mean problem.
Properties of the objective function and solution
Because of the non-linearity of the estimator (and of the $r_{\mathrm{BNM}}$ function) $A(B,\delta)$ is not quasi-convex in $B^{2}$ for the chosen non-linear shrinkage rule. The optimal solution is obtained by solving a minimax problem by using a discrete approximation over the least-favorable prior over $B$ (Chamberlain, 2000).
This differs from our framework where quasi-convexity arises as we restrict the class of estimators to be linear. A formal argument is provided in the proof of Theorem 1. Such quasi-convexity simplifies estimation here, since the choice is not only for the estimator, but also for the class of designs and sample size allocations.